# Local Advice and Local Decompression

### v1

@@ 4. LCLs on Graphs with Subexponential Growth
We consider the nodes in $Z$ in order of their IDs, and we greedily compute a Maximal Independent Set $Z^{\prime}$. Consider the nodes in $Z^{\prime}$ sorted by their IDs. We assign to the $j$th node of $Z^{\prime}$ the $j$th bit of $B$. By Lemma 4, the nodes that received a $1$ for the encoding of the solution of $\Pi$ form an independent set in $G$.
@@
We now show that we can make the ratio between nodes holding a $1$ and those holding a 00 an arbitrarily small constant. There are two places in the schema where we assign $1$s, and those are the following.
  • When we encode the color of the center of the clusters: in this case, the number of $1$s in a cluster $C_{v}$ is upper bounded by $h=20cx+9$, and there are at least $y-h$ nodes in $P_{v}$ holding 00s; by choosing $c$ small enough, we can obtain any desired ratio between the number of $1$s and the number of 00s.
  • When we encode the solution for $\Pi$: in this case, by picking $r$ large enough, we can make the encoded bit-string $B$ arbitrarily small compared to the size of $Z^{\prime}$, that is, the nodes that encode $B$; note that the nodes in $Z^{\prime}$ that do not take part in the encoding of $B$ have a 00 assigned.
5. Balanced Orientations
@@
We start by considering the case in which all nodes have even degree. Note that, in this case, a balanced orientation always exists (though, without advice, the problem of computing such an orientation requires $\Omega(n)$ rounds, e.g., in a cycle). Later, we will show how to extend this result to the more general case. Finally, we will show that that balanced orientations can be used to solve variants of the edge coloring problem. We start by proving the following lemma.
Lemma 0.
Let $\mathcal{G}_{\Delta}$ be the family of graphs of maximum degree $\Delta$, where it is also satisfied that all nodes have an even degree. Let $\Pi_{\Delta}$ be the problem of orienting the edges of a given graph $G\in\mathcal{G}_{\Delta}$ such that each node $v$ satisfies $\deg_{\mathrm{in}}(v)=\deg_{\mathrm{out}}(v)$, where $\deg_{\mathrm{in}}(v)$ and $\deg_{\mathrm{out}}(v)$ are, respectively, the amount of incoming and outgoing edges of $v$. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=\Delta^{O(\alpha)}$.
By applying Lemma 2, which allows us to convert the composable advice schema of Lemma 1 into a fixed-length one, we obtain the following.
Corollary 0.
@@
We devote the rest of the section to prove Lemma 1. Let $G$ be a graph in $\mathcal{G}_{\Delta}$. We start with a simple observation. The balanced orientation problem can be easily solved in constant time if we are given an oracle for consistently orienting cycles. In fact, consider the following algorithm:
  1. (1)
    Starting from $G$, nodes construct a virtual graph $G^{\prime}$ where all nodes have degree $2$, as follows. Each node $v$ of degree $2d$ makes $d$ copies of itself. The copy number $i\in\{1,\ldots,d\}$ is incident to the $(2i-1)$-th and $2i$-th edges of $v$, taken in some arbitrary fixed order (e.g., by sorting the neighbors of $v$ by their IDs). Note that $G^{\prime}$ is a collection of cycles.
  2. (2)
@@
Let $\beta=\gamma_{0}=2$. We show that, for any $\gamma\geq\gamma_{0}$, for any constant $c$, and for any $\alpha\geq\max\{\gamma^{3}\beta/c,\gamma^{3}\beta\}$, there exists a variable-length $(\mathcal{G}_{\Delta},\Pi_{\Delta},\beta,\Delta^{O(\alpha)})$-advice schema satisfying that in each $\alpha$-radius neighborhood the bit-holding nodes are at most $\gamma_{0}$. Since this condition satisfies the requirements of Definition 4, we would obtain Lemma 1.
Let $G^{\prime\prime}$ be the graph obtained from $G^{\prime}$ by removing all cycles of length at most $r$, for some parameter $r$ to be fixed later. We prove that there exists a subset $S^{\prime\prime}$ of the nodes of $G^{\prime\prime}$ satisfying the following properties, where $S$ is defined as the set of nodes $v$ of $G$ satisfying that at least one copy of $v$ is in $S^{\prime\prime}$.
  1. (1)
    All nodes that are part of $G^{\prime\prime}$ are within distance $r$ (in $G^{\prime\prime}$) from a node of $S^{\prime\prime}$.
  2. (2)
    The nodes in $S$ must be at pairwise distance (in $G$) at least $3\alpha$.
  3. (3)
    If two nodes of $G^{\prime\prime}$ are copies of the same node in $G$, then they cannot both be in $S^{\prime\prime}$.
@@
  • The second bit of $v^{\prime\prime}$ is set to $1$ if the cycle is oriented from $v^{\prime\prime}$ to $u^{\prime\prime}$, and to 00 otherwise.
Observe that bit-holding nodes that correspond to different nodes in $S$ are at distance at least $13\alpha-2\geq 2\alpha+\gamma^{3}\beta-2\geq 2\alpha+$ 1, and hence each $\alpha$-radius neighborhood contains at most $2$ bit-holding nodes, as required. A proper orientation for the cycles of $G^{\prime}$ can be recovered as follows.
@@
Thus, we obtain a balanced orientation for $G$ by spending $O(r)$ rounds. Hence, the only remaining thing to be done is to prove the existence of $S^{\prime\prime}$ when $r=\Delta^{O(\alpha)}$. Let $\bar{r}=\varepsilon r$, for a small-enough constant $\varepsilon$ to be fixed later. We start by computing a set $\bar{S}$ of nodes of $G^{\prime\prime}$ that satisfies the following.
@@
  • The distance of any pair of distinct nodes in $\bar{S}$ is at least $12\bar{r}+$ 1.
@@
  • The number of variables that affect the same event is bounded by $\frac{\Delta}{2}\Delta^{3\alpha+1}$. In fact, if we consider a node $v$ of $G$, in its $3\alpha$-radius neighborhood there are at most $\Delta^{3\alpha+1}$ nodes. For each node $u$ among these, there are at most $\Delta/2$ paths that pass through $u$ in $G^{\prime}$. For each path $P$ that passes through $u$, since nodes of $\bar{S}$ are at distance at least $12\bar{r}+$ 1, and each node in $\bar{S}$ is shifted by at most $\bar{r}$, only one node of $\bar{S}$ that lies on $P$ can reach $u$ after being shifted.
We can upper bound $d$ as $\Delta^{3\alpha+1}\bar{r}\cdot\frac{\Delta}{2}\Delta^{3\alpha+1}\leq\bar{r}\Delta^{6\alpha+3}$. Moreover, we can upper bound the probability that a bad event happens on a node $v$ as follows. As mentioned before, the number of nodes that can be shifted into the $3\alpha$-radius neighborhood of $v$ is upper bounded by $\Delta^{3\alpha+1}\Delta/2$. The probability that a shifted node lands in the $3\alpha$-radius neighborhood of a node can be upper bounded by $\frac{\Delta^{3\alpha+1}}{\bar{r}}$, since a shifted node could land in $\bar{r}$ possible nodes, but there are at most $\Delta^{3\alpha+1}$ nodes in the $3\alpha$-radius neighborhood of $v$. A bad event happens on $v$ when at least two nodes are shifted in the $3\alpha$-radius neighborhood of $v$. Hence, by a union bound over all possible pairs, we can upper bound the failure probability on $v$ as $p\leq{\Delta^{3\alpha+1}\Delta/2\choose 2}(\frac{\Delta^{3\alpha+1}}{\bar{r}})^{2}\leq\frac{\Delta^{12\alpha+6}}{\bar{r}^{2}}$. Note that $pd\leq\Delta^{18\alpha+9}/\bar{r}$. Hence, we can pick $r\in\Delta^{O(\alpha)}$ such that $epd\leq e\Delta^{18\alpha+9}/\bar{r}<1$. By the Lovász Local Lemma, this implies that an assignment of shifts satisfying the requirements exists.
@@
We can extend the proof of Lemma 1 to the case in which nodes are allowed to have an odd degree and obtain an orientation in which each node $v$ satisfies $|\deg_{\mathrm{in}}(v)-\deg_{\mathrm{out}}(v)|\leq 1$, as follows. Nodes can compute a virtual graph $G^{\prime}$ that is a collection of paths and cycles, such that each node in $G$ is the endpoint on at most one path passing through it. Then, nodes can compute $G^{\prime\prime}$ by discarding paths that are shorter than $r$. After this, we use the same advice schema as in Lemma 1 to obtain consistently oriented paths. By mapping the orientation of the paths to an orientation of the edges of $G$, we obtain $|\deg_{\mathrm{in}}(v)-\deg_{\mathrm{out}}(v)|\leq 1$, as required. Thus, we obtain the following corollary.
Corollary 0.
Let $\mathcal{G}_{\Delta}$ be the family of graphs of maximum degree $\Delta$. Let $\Pi_{\Delta}$ be the problem of orienting the edges of a given graph $G\in\mathcal{G}_{\Delta}$ such that each node $v$ satisfies $|\deg_{\mathrm{in}}(v)-\deg_{\mathrm{out}}(v)|\leq 1$, where $\deg_{\mathrm{in}}(v)$ and $\deg_{\mathrm{out}}(v)$ are, respectively, the amount of incoming and outgoing edges of $v$. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=\Delta^{O(\alpha)}$.
Corollary 0.
@@
Let $\mathcal{G}_{\Delta}$ be the family of graphs maximum degree $\Delta$, where all nodes have even degree. Let $\Pi$ be the problem of computing an edge coloring of $G$ with $2$ colors, say red and blue, such that, for each node $v$ of degree $d$, $v$ has the same amount of incident red and blue edges. This problem is called splitting. We show that, for the problem $\Pi$, there exists a composable schema.
@@
Extension to $\Delta$-edge coloring on bipartite $\Delta$-regular graphs, when $\Delta$ is a power of $2$.
@@
For each $1\leq i\leq\log\Delta$, we define the problem $\Pi_{i}$ as the problem of computing an edge coloring of $G$ with $2^{i}$ colors, such that each color-class induces a $\Delta/2^{i}$-regular subgraph. Observe that, by assumption, for each $i$, $\Delta/2^{i}$ is an integer. Note that $\Pi_{1}$ is exactly the splitting problem, and solving $\Pi_{i+1}$ given a solution for $\Pi_{i}$ is exactly the problem of computing a splitting for each subgraph given by the solution for $\Pi_{i}$. Observe that $\Pi_{\log\Delta}$ is exactly the problem of computing a $\Delta$-edge coloring. Thus, we can solve $\Pi_{1}$ given advice for splitting, and we can then solve $\Pi_{2}$ given a solution for $\Pi_{1}$ and two separate advices for splitting, one for each subgraph given by $\Pi$. In total, in order to solve $\Pi_{\log\Delta}$, we thus need $O(\Delta)$ advices for splitting. Thus, we obtain the claimed results by applying Lemma 1 with $k=O(\Delta)$ on the schema given by Corollary 5.
6. $\Delta$-coloring of $\Delta$-colorable graphs
Let $G$ be a graph with maximum degree $\Delta$. We show that any such graph that is $\Delta$-colorable can be colored with $\Delta$ colors in $O(\log\Delta\cdot\sqrt{\Delta\log\Delta})$ rounds using $1$ bit of initial advice. More precisely, we design a composable advice schema for the $\Delta$-coloring problem, and then we apply Lemma 2 to obtain a schema that uses $1$ bit per node.
Theorem 1.

### v2

@@ Appendix A LCLs on graphs with sub-exponential growth
We consider the nodes in $Z$ in order of their IDs, and we greedily compute a Maximal Independent Set $Z^{\prime}$. Consider the nodes in $Z^{\prime}$ sorted by their IDs. We assign to the $j$th node of $Z^{\prime}$ the $j$th bit of $B$. By Lemma˜A.5, the nodes that received a $1$ for the encoding of the solution of $\Pi$ form an independent set in $G$.
@@
We now show that we can make the ratio between nodes holding a $1$ and those holding a 0 an arbitrarily small constant. There are two places in the schema where we assign $1$s, and those are the following.
  • When we encode the color of the center of the clusters: in this case, the number of $1$s in a cluster $C_{v}$ is upper bounded by $h=20cx+9$, and there are at least yhy-h nodes in $P_{v}$ holding 0s; by choosing $c$ small enough, we can obtain any desired ratio between the number of $1$s and the number of 0s.
  • When we encode the solution for $\Pi$: in this case, by picking $r$ large enough, we can make the encoded bit-string $B$ arbitrarily small compared to the size of $Z^{\prime}$, that is, the nodes that encode $B$; note that the nodes in $Z^{\prime}$ that do not take part in the encoding of $B$ have a 0 assigned.
Appendix B Balanced orientations
@@
We start by considering the case in which all nodes have even degree. Note that, in this case, a balanced orientation always exists (though, without advice, the problem of computing such an orientation requires $\Omega(n)$ rounds, e.g., in a cycle). Later, we will show how to extend this result to the more general case. Finally, we will show that balanced orientations can be used to solve variants of the edge coloring problem. We start by proving the following lemma.
Lemma B.1.
Let $\mathcal{G}_{\Delta}$ be the family of graphs of maximum degree $\Delta$, where it is also satisfied that all nodes have an even degree. Let $\Pi_{\Delta}$ be the problem of orienting the edges of a given graph $G\in\mathcal{G}_{\Delta}$ such that each node $v$ satisfies $\deg_{\mathrm{in}}(v)=\deg_{\mathrm{out}}(v)$, where $\deg_{\mathrm{in}}(v)$ and $\deg_{\mathrm{out}}(v)$ are, respectively, the number of incoming and outgoing edges of $v$. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=\Delta^{O(\alpha)}$.
By applying Lemma˜E.3, which allows us to convert the composable advice schema of Lemma˜B.1 into a fixed-length one, we obtain the following.
Corollary B.2.
@@
We devote the rest of the section to prove Lemma˜B.1. Let $G$ be a graph in $\mathcal{G}_{\Delta}$. We start with a simple observation. The balanced orientation problem can be easily solved in constant time if we are given an oracle for consistently orienting cycles. In fact, consider the following algorithm:
  1. 1.
    Starting from $G$, nodes construct a virtual graph $G^{\prime}$ where all nodes have degree $2$, as follows. Each node $v$ of degree $2d$ makes $d$ copies of itself. The copy number $i\in\{1,\ldots,d\}$ is incident to the (2i1)(2i-1)-th and $2i$-th edges of $v$, taken in some arbitrary fixed order (e.g., by sorting the neighbors of $v$ by their IDs). Note that $G^{\prime}$ is a collection of cycles.
  2. 2.
@@
Let $\beta=\gamma_{0}=2$. We show that, for any $\gamma\geq\gamma_{0}$, for any constant $c$, and for any $\alpha\geq\max\{\gamma^{3}\beta/c,\gamma^{3}\beta\}$, there exists a variable-length $(\mathcal{G}_{\Delta},\Pi_{\Delta},\beta,\Delta^{O(\alpha)})$-advice schema satisfying that in each $\alpha$-radius neighborhood there are at most $\gamma_{0}$ bit-holding nodes. Since this condition satisfies the requirements of Definition˜3.4, we would obtain Lemma˜B.1.
Let $G^{\prime\prime}$ be the graph obtained from $G^{\prime}$ by removing all cycles of length at most $r$, for some parameter $r$ to be fixed later. Let $f:V(G^{\prime\prime})\rightarrow V(G)$ be the function defined as follows: for each node $v$ in $G^{\prime\prime}$ that is a copy of node $u$ in $G$, $f$ maps $v$ to $u$. We prove that there exists a subset $S^{\prime\prime}$ of the nodes of $G^{\prime\prime}$ satisfying the following properties, where $S:=\{f(v)\mid v\in S^{\prime\prime}\}$.
  1. 1.
    $S^{\prime\prime}$ is an $r$-dominating set of $G^{\prime\prime}$, that is, all nodes that are part of $G^{\prime\prime}$ are within distance $r$ (in $G^{\prime\prime}$) from a node of $S^{\prime\prime}$.
  2. 2.
    If $v^{\prime\prime},u^{\prime\prime}\in S^{\prime\prime}$, then $\operatorname{dist}_{G}(f(v^{\prime\prime}),f(u^{\prime\prime}))\geq 3\alpha$.
@@
  • The second bit of $v^{\prime\prime}$ is set to $1$ if the cycle is oriented from $v^{\prime\prime}$ to $u^{\prime\prime}$, and to 0 otherwise.
Observe that bit-holding nodes that correspond to different nodes in $S$ are at distance at least $3\alpha-2\geq 2\alpha+\gamma^{3}\beta-2\geq 2\alpha+1$, and hence each $\alpha$-radius neighborhood contains at most $2$ bit-holding nodes, as required. A proper orientation for the cycles of $G^{\prime}$ can be recovered as follows.
@@
Thus, we obtain a balanced orientation for $G$ by spending $O(r)$ rounds. An example is depicted in Figure˜1.
(a)Refer to caption(b)Refer to caption(c)Refer to caption
Figure 1: (a) The original graph $G$. (b) The graph $G^{\prime}$, obtained by decomposing $G$ into cycles. By removing the three gray cycles (solid, dashed, and dotted), which are considered short and can be oriented without advice, we obtain $G^{\prime\prime}$. (c) For each directed edge, exactly one node belongs to $S^{\prime\prime}$. Such nodes receive advice encoding the orientation of these edges. We use the Lovász Local Lemma to prove that it is possible to choose $S^{\prime\prime}$ such that their corresponding images in $G$ are far enough from each other, and such that all nodes can recover the orientation of all the cycles they belong to by inspecting $G$ up to a bounded distance.
The only remaining thing to be done is to prove the existence of $S^{\prime\prime}$ when $r=\Delta^{O(\alpha)}$. Let $\bar{r}=\varepsilon r$, for a small-enough constant $\varepsilon$ to be fixed later. We start by computing a set $\bar{S}$ of nodes of $G^{\prime\prime}$ that satisfies the following.
@@
  • The distance of any pair of distinct nodes in $\bar{S}$ is at least $2\bar{r}+1$.
@@
  • The number of variables that affect the same event is bounded by $\frac{\Delta}{2}\Delta^{3\alpha+1}$. In fact, if we consider a node $v$ of $G$, in its $3\alpha$-radius neighborhood there are at most $\Delta^{3\alpha+1}$ nodes. For each node $u$ among these, there are at most $\Delta/2$ paths that pass through $u$ in $G^{\prime}$. For each path $P$ that passes through $u$, since nodes of $\bar{S}$ are at distance at least $2\bar{r}+1$, and each node in $\bar{S}$ is shifted by at most $\bar{r}$, only one node of $\bar{S}$ that lies on $P$ can reach $u$ after being shifted.
We can upper bound $d$ as $\Delta^{3\alpha+1}\bar{r}\cdot\frac{\Delta}{2}\Delta^{3\alpha+1}\leq\bar{r}\Delta^{6\alpha+3}$. Moreover, we can upper bound the probability that a bad event happens on a node $v$ as follows. As mentioned before, the number of nodes that can be shifted into the $3\alpha$-radius neighborhood of $v$ is upper bounded by $\Delta^{3\alpha+1}\Delta/2$. The probability that a shifted node lands in the $3\alpha$-radius neighborhood of a node can be upper bounded by $\frac{\Delta^{3\alpha+1}}{\bar{r}}$, since a shifted node could land in $\bar{r}$ possible nodes, but there are at most $\Delta^{3\alpha+1}$ nodes in the $3\alpha$-radius neighborhood of $v$. A bad event happens on $v$ when at least two nodes are shifted in the $3\alpha$-radius neighborhood of $v$. Hence, by a union bound over all possible pairs, we can upper bound the failure probability on $v$ as
$p\leq\binom{\Delta^{3\alpha+1}\Delta/2}{2}\left(\frac{\Delta^{3\alpha+1}}{\bar{r}}\right)^{2}\leq\frac{\Delta^{12\alpha+6}}{\bar{r}^{2}}.$
Note that $pd\leq\Delta^{18\alpha+9}/\bar{r}$. Hence, we can pick $r\in\Delta^{O(\alpha)}$ such that $epd\leq e\Delta^{18\alpha+9}/\bar{r}<1$. By the Lovász Local Lemma, this implies that an assignment of shifts satisfying the requirements exists.
@@
We can extend the proof of Lemma˜B.1 to the case in which nodes are allowed to have an odd degree and obtain an orientation in which each node $v$ satisfies $|\deg_{\mathrm{in}}(v)-\deg_{\mathrm{out}}(v)|\leq 1$, as follows. Nodes can compute a virtual graph $G^{\prime}$ that is a collection of paths and cycles, such that each node in $G$ is the endpoint on at most one path passing through it. Then, nodes can compute $G^{\prime\prime}$ by discarding paths that are shorter than $r$. After this, we use the same advice schema as in Lemma˜B.1 to obtain consistently oriented paths. By mapping the orientation of the paths to an orientation of the edges of $G$, we obtain $|\deg_{\mathrm{in}}(v)-\deg_{\mathrm{out}}(v)|\leq 1$, as required. Thus, we obtain the following corollary.
Corollary B.3.
Let $\mathcal{G}_{\Delta}$ be the family of graphs of maximum degree $\Delta$. Let $\Pi_{\Delta}$ be the problem of orienting the edges of a given graph $G\in\mathcal{G}_{\Delta}$ such that each node $v$ satisfies $|\deg_{\mathrm{in}}(v)-\deg_{\mathrm{out}}(v)|\leq 1$, where $\deg_{\mathrm{in}}(v)$ and $\deg_{\mathrm{out}}(v)$ are, respectively, the number of incoming and outgoing edges of $v$. Then, there exists a $(\mathcal{G}_{\Delta},\Pi_{\Delta},\gamma_{0},A,T)$-composable advice schema, where $\gamma_{0}=O(1)$, $A(c,\gamma)=\Theta(\gamma^{3})$, and $T(\alpha,\Delta)=\Delta^{O(\alpha)}$.
Corollary B.4.
@@
Let $\mathcal{G}_{\Delta}$ be the family of graphs maximum degree $\Delta$, where all nodes have even degree. Let $\Pi$ be the problem of computing an edge coloring of $G$ with $2$ colors, say red and blue, such that, for each node $v$ of degree $d$, $v$ has the same number of incident red and blue edges. This problem is called splitting. We show that, for the problem $\Pi$, there exists a composable schema.
@@
[…]
@@
For each $1\leq i\leq\log\Delta$, we define the problem $\Pi_{i}$ as the problem of computing an edge coloring of $G$ with $2^{i}$ colors, such that each color-class induces a $\Delta/2^{i}$-regular subgraph. Observe that, by assumption, for each $i$, $\Delta/2^{i}$ is an integer. Note that $\Pi_{1}$ is exactly the splitting problem, and solving $\Pi_{i+1}$ given a solution for $\Pi_{i}$ is exactly the problem of computing a splitting for each subgraph given by the solution for $\Pi_{i}$. Observe that $\Pi_{\log\Delta}$ is exactly the problem of computing a $\Delta$-edge coloring. Thus, we can solve $\Pi_{1}$ given advice for splitting, and we can then solve $\Pi_{2}$ given a solution for $\Pi_{1}$ and two separate advices for splitting, one for each subgraph given by $\Pi$. In total, in order to solve $\Pi_{\log\Delta}$, we thus need $O(\Delta)$ advices for splitting. Thus, we obtain the claimed results by applying Lemma˜E.1 with $k=O(\Delta)$ on the schema given by Corollary˜B.5.
Appendix C $\Delta$-coloring of $\Delta$-colorable graphs
Let $G$ be a graph with maximum degree $\Delta$. We show that any such graph that is $\Delta$-colorable can be colored with $\Delta$ colors in $O(\log\Delta\cdot\sqrt{\Delta\log\Delta})$ rounds using $1$ bit of initial advice. More precisely, we design a composable advice schema for the $\Delta$-coloring problem, and then we apply Lemma˜E.3 to obtain a schema that uses $1$ bit per node.
Theorem C.1.